### v1

# TIAM - A Metric for Evaluating Alignment in Text-to-Image Generation

Abstract
The progress in the generation of synthetic images has made it crucial to assess their quality. While several metrics have been proposed to assess the rendering of images, it is crucial for Text-to-Image (T2I) models, which generate images based on a prompt, to consider additional aspects such as to which extent the generated image matches the important content of the prompt. Moreover, although the generated images usually result from a random starting point, the influence of this one is generally not considered. In this article, we propose a new metric based on prompt templates to study the alignment between the content specified in the prompt and the corresponding generated images. It allows us to better characterize the alignment in terms of the type of the specified objects, their number, and their color. We conducted a study on several recent T2I models about various aspects. An additional interesting result we obtained with our approach is that image quality can vary drastically depending on the latent noise used as a seed for the images. We also quantify the influence of the number of concepts in the prompt, their order as well as their (color) attributes. Finally, our method allows us to identify some latent seeds that produce better images than others, opening novel directions of research on this understudied topic.
1 Introduction
The ability to generate synthetic images with neural models made significant advancements from the advent of the first GANs [7, 20]. More recently diffusion-based models  [11, 24, 25, 17, 1] have further pushed the boundaries of image synthesis by progressively denoising an initial latent noise to generate high-quality images. In parallel to these progresses, the question of evaluating the quality of the generated synthetic images has always been a delicate issue and has become a research question in itself. To address this problem [28], several metrics were proposed [8, 26, 32] but they suffer from various limits [2, 4]. The most recent models are conditioned on textual image descriptions, allowing fine control of the output. It nevertheless added a challenge to evaluate their outputs, namely to estimate to which extent the synthetic image generated corresponds to the textual description it was conditioned on.
Firstly, although Text-to-Image (T2I) models demonstrate strong semantic and compositional capabilities, achieving a visually pleasing image that aligns with the desired condition often requires the generation of multiple images to obtain a suitable one. We propose that a reliable generative model should exhibit alignment with the condition specified in the prompt, irrespective of the initial latent noise. To address and study the variability of results, we introduce a novel metric to assess the success rate of generative models according to a prompt, Text-Image Alignment Metric [^1] Source code: https://github.com/grimalPaul/TIAM abbreviated as TIAM. The initial latent noise plays a crucial role in our metric, enabling us to investigate its impact. We have demonstrated that certain initial latent noise configurations outperform others. This finding suggests the possibility of selecting more effective initial latent noise to generate desired content.
Secondly, recent research efforts [5, 22, 25, 27] have shown that text-conditioned diffusion models suffer from three main issues relating to the text-image alignment, that is the alignment between the expected content expressed in the textual prompt and the content that is actually contained in the generated image: (i) catastrophic neglect where one or more elements of the prompt are not generated or sometimes mixed, (ii) attribute binding where attributes are bound to the wrong entities and (iii) attribute leaking where attributes specified in the prompt are correctly bound but some other elements in the scene are also wrongly bound with this attribute (Fig 1).
With TIAM, we propose to analyze the success rate of generative models under the scope of catastrophic and attribute-binding issues. For the latter, we propose a solid method to evaluate colors aligned with human perception.
Refer to caption Refer to caption Refer to caption
(a) Catastrophic (b) Attribute (c) Attribute
neglect binding leaking
Figure 1: Respectively, pictures generated with the prompts “a photo of a lion and a bear”, “a photo of a blue cat and a yellow car”, and “a photo of a red bus driving down the street” generated with Stable diffusion v1.4. (a) The bear is missing, (b) the attributes are swapped (c) the wall becomes red because of the adjective that characterizes the bus.
Finally, the investigation of the influence of some words and attributes remains largely understudied. Tang et al. [27] provided some insights by examining the impact of words on generated outcomes. Using TIAM, we provide further insights into the relationship between textual conditioning and generative results. Relying on prompt template rather than natural prompts, TIAM allows quantifying the prompt-image alignment with regard to syntactic aspects, in particular the importance of the position of the main entities in the prompt.
In summary, our main contributions are the following: (i) We present a new metric based on prompt templates to quantify automatically the performances of T2I models in terms of prompt-image alignment (ii) We conduct an in-depth study of several diffusion models and quantify their behavior relating to catastrophic neglect and attribute binding issues (iii) We study the influence of the initial diffusion noise (seed) for all these models. The main insights resulting from this study are (a) the alignment performances of most T2I models drop significantly with the number of objects specified in the prompt (b) in practice there exists some seed that systematically leads to better results; TIAM allows us to identify them and they still provide better results with objects that are out of our study domain (c) most T2I models succeed in attributing color to one object but the performance drop even further with more objects.
2 Related Work
Text-to-image models
Among the current state-of-the-art models, diffusion models [11] have demonstrated remarkable performance. These models introduce noise to images and learn to denoise the added degradation. During inference, the models iteratively denoise a noise sampling from a Gaussian distribution, resulting in a reconstructed image. The utilization of free classifier guidance [12] allows users to express their thoughts by writing a prompt using natural language. The diffusion process is then guided by the encoded text representation, leveraging foundational models like CLIP [19] or T5 [21]. It is important to note that the results are highly stochastic, often requiring the generation of multiple images with different Gaussian noise inputs to align with the user’s desired texts and preferences. Among the well-known models in this field, notable mentions include Imagen [25], Dall-E 2 [22], Latent Diffusion Model (LDM) [24] and Ediff-I [1].
T2I Evaluation
The evaluation of generated images often relies on metrics such as Inception Score (IS) [26] and Fréchet Inception Distance (FID) [8], which are commonly used to assess their quality and fidelity. However, these metrics do not capture the alignment between the provided condition(s) and the generated images. To address this issue, several methods have been proposed to measure whether the content of the generated images reflects the given conditions. One approach involves utilizing classification models [23] and object detection models (Semantic Object Accuracy, abbreviated as SOA) [10] to determine whether the generated images contain the specified objects or adhere to certain criteria. SOA uses actual prompts from COCO while our approach relies on prompt templates that allows a finer analysis of the influence of each element in the prompt. Another method specific to text-image alignment involves employing visual score similarity measures. It leverages contrastive models such as CLIP, which compute a similarity score between representations of images and text. Although CLIP has a great average semantic representation, it has poor compositional understanding [31], limiting a fine evaluation of text-image alignment. Our metric addresses this issue by leveraging a high-performance detector and controlling the requested attributes in an explainable manner and aligning with human perception.
Some benchmarks such as DrawBench [25] propose a limited set of text prompts covering 11 skills to evaluate models. To gain a more comprehensive understanding of the performance of T2I models, we approach the issue by focusing on catastrophic neglect and attribute binding. To overcome results that are very sensitive to asked objects in prompt and evaluate the model’s performance accurately, we propose a methodology that tests all combinations of objects and attributes and generates multiple images. By testing all the combinations in the prompts, we can assess the model’s success rate in capturing and representing the desired attributes. This approach offers a comprehensive evaluation, free from the limitations of specific labels, resulting in a more precise and less biased assessment.
3 Method
We propose a new method to measure the success rate of generative models covering catastrophic neglect and/or attribute binding. First, we generate multiple prompts given a set of word labels and possibly attributes. We then generate multiple images for each prompt and detect if the expected asked elements are detected on the image. Finally, we compute the final score.
3.1 TIAM Text-Image Alignment Metric
Our approach is based on templates that are used to generate sets of prompts. We adopt a formalism inspired by the disentangled representation theory [9, 29]. The prompts contain $N$ objects, each of which can be qualified by an attribute. Hence the object at position $i$ is a token belonging to the set $\mathcal{O}_{i}$ and qualified by an attribute in the set $\mathcal{A}_{i}$. For example, let us consider the template “a photo of det($o_{1}$, $a_{1}$) $a_{1}$ $o_{1}$ and det($o_{2}$, $a_{2}$) $a_{2}$ $o_{2}$” where $o_{i}\in\mathcal{O}_{i}$, $a_{i}\in\mathcal{A}_{i}$, and det($o_{i}$, $a_{i}$) is a determinant that depends on the object or attribute if present. If $\mathcal{O}_{1}=\mathcal{O}_{2}=${‘car’,‘bike’,‘truck’} and $\mathcal{A}_{1}=\mathcal{A}_{2}=${‘blue’,‘green’,‘red’}, then it can produce prompts such as “a photo of a blue car and a red truck” or “a photo of a green bike and a blue car”.
For such a template $t$, the generic expression of the text-image alignment metric (TIAM) is defined as:
$\underset{\begin{subarray}{c}\chi\sim\mathcal{N}(0,I)\\ z\in\mathcal{Z}\end{subarray}}{\mathbb{E}}\left[f\left(G(\chi,t(z)),y(z)\right)\right]\quad\text{with:}\quad\mathcal{Z}=\prod_{i=1}^{N}\left(\mathcal{A}_{i}\times\mathcal{O}_{i}\right)$ (1)
where the prompt is instanced from the template $t$ and its “latent concepts” $z\in\mathcal{Z}$, $y(z)$ is the labels that relate to the expected content of the synthetic image generated by the model $G()$ (conditioned by the prompt) from a seed (using to generate a latent noise) $\chi$. $f$ is a scoring function that compares the ground truth $y(z)$ to the output of a model that detects objects or produces segmentation maps from the synthetic image, the resulting score being in $\{0,1\}$ depending on whether the content matches the ground truth or not (see subsection 3.3).
With the definition of Equation 1, a template can generate $\sum_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$ different prompts, where |.||.|| . | is the cardinal of the set. However, we restrict in practice the inference of the template such that an object or an attribute can not appear twice in the prompt. In a simple case where the attributes and the objects set are the same at each position, the number of unique prompts is given by:
Proposition 1: if $|\mathcal{O}|\geq N$, $|\mathcal{A}|\geq N$ and $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$ and $\forall(i,j)\in[\![1,N]\!]^{2}$ s.t $i<j$, we force $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$, thus the number of unique prompts generated in the context of Equation 1 is $\frac{|\mathcal{O}|!|\mathcal{A}|!}{\left(|\mathcal{O}|-N\right)!\left(|\mathcal{A}|-N\right)!}$
The proof is given in the supplementary material A. We also derive the more general case with possibly different attributes and objects at each position. The global evaluation process is summarized in Fig 2.
Figure 2: Overview of the evaluation pipeline. (1) Generate a dataset of prompts (2) Generate $n\geq 16$ images per prompt (3) Detect if the requested labels are present in the image (4) Compute TIAM. In this example, we do not define attributes.
3.2 Attribute definition
We study the color attributes in this paper, but this metric could be applied under the scope of another type of attribute (e.g., size or texture). Selecting colors that are aligned with human perception is not trivial because of the infinity of possibilities. We base our choice on the work of Berlin and Kay [3]. They define eleven universal basic colors $\mathcal{C}$: white, black, red, green, blue, brown, purple, pink, orange, yellow, and grey. They asked individuals to select an array of 329 colors (provided by the Munsell Colors Company), the chips that correspond to each basic color, and to select the most typical examples. We use the results of the most typical examples of American English and convert the colors from the Munsell System to the CIELAB space. This one offers the advantage of a color distribution more aligned with human perception and a suitable space to compare color differences.
We remark that brown and orange were respectively too close to black and white and we remove them (more information provided in supplementary material C. We set
[…]
However, we use the following set as reference colors (to compare with $\mathcal{A}$): white, black, red, green, blue, purple, pink, yellow.
To define if $\mathcal{G()}$ correctly assigns $a_{i}$ and $o_{i}$, we use a segmentation model that delimits $o_{i}$ if present on the image. We compare all the pixels inside the segmentation maps with our reference colors in the CIELAB space. If we detect at least a proportion of 40% of $a_{i}$, we consider a successful binding.
3.3 Main implementation details
The template $t$ always includes the start mention “a photo of”. Indeed, in our case, the detector is trained on real images and we suppose that it would therefore have more difficulty managing e.g. sketches or 3D renders. The general form of the templates is that given as an example above, by varying the syntactic context. For the detection and segmentation tasks we use YOLOv8 [13], a state-of-the-art model pre-trained on the 80 COCO labels [15]. We fixed the object confidence threshold to 0.25 but conducted a study asserting that a higher threshold (0.4 to 0.8) does not change the relative order performance of the models. The results can be found in the supplementary material F, where we provide a detailed description of the labels used in each experiment and their combination within the generated prompts.
The scoring function $f$ mainly focuses on true positives, considering the alignment as successful if we find at least one of each named object in the prompt. If the objects are characterized by an attribute, each object must be present with its correct attribute at least once in the image to be marked as a success. The criterion is therefore strict on the alignment aspect, requiring the presence of all the named objects and attributes. This approach differs from CLIP scores, which provide a general trend without reflecting a comprehensive explanation of the results. It nevertheless does not penalize the presence of other objects (false positive) since it is not prohibitive in a generative context. $f$ prevent also element mixing. Indeed, if the segmentation masks of two different objects overlap with an IoU value of 0.95 or higher, we remove both objects before calculating the score with the remaining detected objects.
4 Study
We conduct an in-depth study to characterize and quantify the limits of several diffusion models, which differ by the text-conditioning method, the architecture, or the inference process. We consider Stable Diffusion v1-4 (SD 1.4) [^2] https://huggingface.co/CompVis/stable-diffusion-v1-4 and Stable Diffusion v2 (SD 2) [^3] https://huggingface.co/stabilityai/stable-diffusion-2-base based on the LDM model [24] that uses a fixed pretrained text encoder CLIP [19], which is based on an auto-regressive architecture. Both versions are trained with the same variational autoencoder (VAE), but two different U-nets (the part that learns to denoise image). The CLIP models that guide the U-nets also differ between SD 1.4 and SD 2. In addition, we evaluate the two SD models with Attend-and-Excite (A&E) [5], an optimized inference process using the cross-attention maps to attend the subject tokens in the prompt. We follow the author’s implementation and only attend to the token of object $o_{i}$, even if an attribute characterizes the object. We also consider an unCLIP model [22, 14] conditioned with a CLIP image prior and a CLIP text embedding. Finally, we study DeepFloyd IF (IF) [^4] https://github.com/deep-floyd/IF a cascade diffusion inspired by [25] conditioned using a T5 XXL [21] text encoder. Other implementation details are reported in the supplementary material G.
4.1 Preliminary: performance drop with 2 objects
We report a first experiment to illustrate the general framework of our study.
We consider various labels from COCO (the exhaustive list is in supplementary material B), leading to a set of $|\mathcal{O}|=24$ object labels. For all possible couples of labels $(o_{i},o_{j})$ with $o_{i}\neq o_{j}$, we make the prompt “a photo of $o_{i}$ and $o_{j}$” (managing the determinant as expected) and generate 64 synthetic images by changing the random seed $\chi$ and estimate the alignment of the prompt with each image with TIAM. As a reference, we also conduct the same experiment with 64 images generated from the simpler prompt “a photo of $o_{i}$”.
We report the result in Tab 1 for all the models considered. The models consistently succeed in generating images with a single object, with a score above 0.95. However, they struggle to generate simultaneously two objects with at most 66% of images correctly generated, while the prompts are minimally simple. In the recent literature in the field, the T2I models usually exhibit a good quantitative score of rendering, e.g. in terms of inception score or FID, which our experience does not call into question. In the vein of Hinz et al. [10] this experiment only shows that there is a problem with alignment between the content of the generated image and the prompt. More precisely than [10], our experiment specifically quantifies to which extent this drop in alignment performance is due to the presence of multiple concepts in the prompt. For the prompts with two objects, we report in Tab 2 the number of times the objects are actually accurately generated with regard to their position in the prompt (the score in Tab 1 requires the presence of both objects to be valid). We observe a tendency for the first $o_{1}$ to be more prevalent than the second $o_{2}$ across all models, a phenomenon that is further investigated in section 4.3.
By comparing the performances of the models, we notice that A&E inference globally improves the score with a large margin for SD v1.4. From Tab 2, it seems that the overperformance of A&E is due to its ability to enhance the occurrence rate of $o_{2}$ by enforcing the minimum excitation of cross-attention maps.
Model 1 object 2 objects
SD 1.4 0.98 0.41
SD 1.4 A&E 0.96 0.64
SD 2 0.99 0.61
SD 2 A&E 0.98 0.65
unCLIP 0.95 0.50
IF 0.99 0.62
Table 1: TIAM according to the number of objects in the prompt, for all 6 generative models considered in the study.
Model $o_{1}$ $o_{2}$
SD v1.4 0.80 0.60
SD 1.4 A&E 0.85 0.75
SD 2 0.83 0.78
SD 2 A&E 0.84 0.80
unCLIP 0.77 0.71
IF 0.86 0.76
Table 2: Proportion of appearance per order in the prompt. $o_{1}$ and $o_{2}$ refer to the position in the template.
4.2 Importance of the random seed
We consider the same 24 labels as in subsection 4.1 and the prompt “a photo of $o_{i}$ and $o_{j}$”. We also generate 64 synthetic images by changing the random seed $\chi$ and compute $f\left(G(\chi,t(z)),y(z)\right)$ for each prompt and corresponding image. However, this time, the 64 seeds are the same for all the $(o_{i},o_{j})$ [^5] Actually, it was the case in subsection 4.1 for the sake of comparison and coherency and we aggregate the performance per seed, for all possible prompts and images, leading to $24\times 23=552$ estimations of text-image alignment per seed.
The results in Fig 3 show that, for all 6 models, the performance varies a lot with regard to the random seed. These results, which to the best of our knowledge have never been identified in the literature of the T2I generative models, may seem surprising. With such models, one usually expects that all seeds have the same chance to generate the specified elements. Actually, the recent tendency rather consists of engineering finely the prompt [18] to optimize the output of T2I models. On the opposite of this tendency, our method opens up new perspectives toward a complementary optimization based on the choice of “performing seeds” (see subsection 4.5).
Overall, there exists a significant disparity between the best and worst seeds. When considering the interquartile ranges and the min-max range through the box-plots of Fig 3, it appears that the difference in average performance between models is much less significant than their own inner difference due to the variation of the seeds.
Obviously, enhancing a dependence “to the seed” is convenient from a practical point of view, but wrong in all strictness since the seeds are drawn from a Gaussian noise in inference. Being more rigorous requires reminding the diffusion models training process [11]. It consists in learning the reverse process of a fixed Markov chain of length $T$ with models that can be interpreted as an equally weighted sequence of denoising autoencoders. These last are trained to denoise their input, considered as a noisy version of an input training image $I$. In the case of latent diffusion models [24], the process is embedded into the latent space of a VAE (encoder $\mathcal{E}$ + decoder $\mathcal{D}$), such that the autoencoders are U-net networks $\epsilon_{\theta}(x_{t},t)$, $t=1\dots T$ that denoise the latent code $x_{t}$ by minimizing the loss $\mathbb{E}_{\mathcal{E}(I),\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}\right]$. During inference, the process starts from a latent code $\chi_{T}\sim\mathcal{N}(0,1)$, denoises it with the U-net $\epsilon_{\theta}$ to get the final latent code $\chi_{0}$, then obtains the synthetic image with VAE decoder as $\mathcal{D}(\chi_{0})$. This last is thus not just dependent on the random seed $\chi_{T}$ but on the U-net parameter $\theta$ as well. These parameters are learned to denoise the latent codes at any step of the diffusion since the reparametrization trick [11] allows to express the final code directly as:
$x_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha_{t}}}\epsilon}{\sqrt{\bar{\alpha_{t}}}}$ (3)
where $\epsilon\sim\mathcal{N}(0,1)$ and the $\alpha_{t}$ depends on the variance of the diffusion process noise (see [11] and the appendix B of [24] for details). Ideally, $\bar{\alpha_{T}}$ is close to one, such that a synthetic image should result from an almost perfect Gaussian noise. However, in practice the process is not perfect, thus the final $\chi_{0}$ obtained for a given $\chi_{T}\sim\mathcal{N}(0,1)$ at inference time is only an approximation of the “ideal $\chi_{0}$” that one could expect with a perfect optimization [6]. Moreover, since each model is trained independently, its latent space is structured in the same vein and the “path” of the $\chi_{t}$ in the latent space may strongly differ from one model to another. Hence, starting from the same $\chi_{T}\sim\mathcal{N}(0,1)$, they result in quite different $\chi_{0}$ while trying to tend to an ideal one. As a consequence, the “good” and “bad” seeds are specific to each model.
Figure 3: TIAM aggregate per seed for 64 seeds. We show that some started latent noise tends to not be converted to an image with two entities regardless of the entities. “+” shows the mean.
The results are contingent upon the seed, indicating that based on the prompt’s specifications, it appears possible to identify initial noises that are more likely to exhibit multiple objects. It highlights the ongoing need for further advancements in reducing reliance on latent variables. To obtain a robust score and avoid the possible impact of the high variability of the seed’s success rate, we define a minimum of images to generate per prompt. We establish it is adequate to generate 32 images per prompt to obtain a robust score. The corresponding experiment is reported in the supplementary material E.
4.3 Catastrophic Neglect
This section examines the behavior of the models as the number of object sets increases and the role of the order of objects in the prompt.
[…]
When examining the occurrence of different objects, similar trends can be observed as in the preliminary experiment (Tab 2). Specifically, the initial objects in the template tend to appear more frequently than objects inserted subsequently. The results for the template containing four objects are presented in Fig 5 (results for two and three objects in supplementary material D). This reinforces the observation that the concept that is expressed earlier in the prompt has more chances to appear in the final image.
For SD, conditioned by CLIP text-encoders, we hypothesize that the decreasing trend in the occurrence of objects as their position becomes more distant in the prompt is partly due to the auto-regressive nature of the encoder. During self-attention, tokens only receive context from the elements to their left (beginning of the prompt). As a result, tokens are devoid of the contexts of subsequent objects, while the following objects carry the context of the earlier words. This leads to a more ambiguous signal from distant tokens for U-Net models.
The explanation is nevertheless less clear with the T5 encoder, where all words have access to each other during self-attention. Used in IF, one can see that the third and fourth words have the same chance to appear, but it is significantly less than the object in the second position, itself below the first concept expressed in the prompt.
We hypothesize that during cross-attention, the models have learned to give more importance to tokens with earlier positions due to the training data, which typically places important elements related to the image at the beginning of the caption. Asserting this explanation would nevertheless require significant work to analyze the original training dataset used to pre-trained the six generative models, which is out of the scope of this paper.
Figure 4: TIAM with 1 to 4 objects per prompt.
Figure 5: The proportion of occurrences of each object, based on its positions in the prompt. The template of the prompt count 4 objects.
Figure 6: MDS on the objects score dissimilarity for SD 1.4.
Finally, we investigated the impact of semantic relationships between objects within the prompt on the ability of the model to generate both of them. Considering a template with two objects, our hypothesis was that the T2I model would fail more often to represent on the same image two objects that are semantically linked. We considered 28 COCO labels from three macro-classes (vehicles, animals, and foods) and generate images using a template with 2 objects. From the resulting TIAM scores, we derived a dissimilarity metric between all objects (see supplementary material H) and project all the labels with Multidimensional Scaling (MDS). The resulting projections (Fig. 18 for SD 1.4) can be interpreted such that the closer two labels are, the more challenging it becomes for the model to represent them together. Across all models, we observe labels from the same class being clustered together, particularly for animals and vehicles. This suggests the presence of semantic proximity. We measured the semantic distance between both named objects with various methods (such as Wu-Palmer, Cosine similarity of CLIP text embeddings or even attention’s key-value representation) for SD 1.4 and SD 2 but the correlation with the TIAM score was only slightly negative. This indicates that the semantic linking has probably either a small or indirect link to the alignment performance, but in any case, further researches need to be conducted to clarify this point.
4.4 Attribute binding
[…]
We start by computing TIAM regardless of the correctness of the attribute and report the results in Tab 3. Without attributes (✗) the scores are close to those of Tab 1 and adding attributes (✓) has a limited impact when one object only is requested. However, it significantly impacts the performance when two objects are present in the prompt.
Figure 7: TIAM computed with object and color ground truth, for one and two colored objects per prompt.
Figure 8: TIAM per color and per object.
1 object 2 objects
Attribute
SD 1-4 0.97 0.95 0.40 0.18
SD 1.4 A&E 0.93 0.94 0.64 0.42
SD 2 0.97 0.97 0.61 0.26
SD 2 A&E 0.96 0.96 0.66 0.32
unCLIP 0.93 0.89 0.61 0.36
IF 0.98 0.90 0.61 0.49
Table 3: TIAM computed with object(s) ground truth only, with prompts containing attributes or not.
When TIAM is estimated by taking into account both the object and color ground truth, we observe that the models fail to assign the appropriate colors to the objects, even when these last are present in the image (Fig. 7). For instance, in the case of SD 1.4, objects are detected correctly in 95% of single-object prompts, but only approximately 45% of them have the requested colors.
In Fig. 8, we report the attribution scores per color and per object to analyze deeper the attributing ability of the models. We see that models fail to generalize the binding to objects when it is uncommon to observe them in that particular color. As expected, it is indeed easier to assign colors to cars and refrigerators compared to animals. The models have likely been trained on photos of cars and refrigerators in various colors during training as it is more common to come across a green car than a green giraffe.
Following the same approach as in section 4.3, we also analyze the results in relation to the position $i$ of attribute $a_{i}$ and object $o_{i}$ in the prompt involving 2 objects. In that case, the models succeed by a large margin to generate and bind the first object. However, knowing that the first object is more often generated we compute a binding success rate, which is the score of correctly attributed objects among the correctly generated objects (Fig. 9). The $o_{2}$ objects, however, continue to be less attributed. This reinforces the finding that the first object in the prompt has a greater influence on the final generation.
Figure 9: Success rate of color attribution w.r.t the detected objects.
In the supplementary material I, we report the results of the binding success rate differentiated by colors for attributes in the first position and attributes in the second position. We observed that the models face greater difficulty in assigning green and blue colors when two objects are involved (parallel with a single label case). It is worth noting that IF performs better than other models.
4.5 Toward noise mining?
To highlight the importance of noise performance, we selected seeds based on their TIAM score. In Fig. 10 we present qualitative results using some of the worst and best seeds found with our approach, and prompts with objects and attributes that were not considered in our study.
Both with SD 2 and IF, the two objects are better represented in the image resulting from the “good” seed, showing that our method has the ability to find seeds that generalize to objects out of the domain of objects used to determine the best seeds.
Figure 10: Generation of images with the same prompt and a “bad” (left) of a “good” (right) random seed. top: SD 2 “a sketch of a tiger and a surfboard, 4k, 8k, ghibli” seed 23, 11. bottom: IF “a photo of space tiger and a rocket” seed 41, 9.
We conducted the same seed selection process for both our worst and best seeds using colors in the prompt (Fig. 11). Once again the images resulting from the best seeds we identified with TIAM better reflect the prompt. With the “bad” seed, SD 2 suffers both from attribute leaking or binding (the “blue moon” is yellow, and the “red lion” is blue), and for both models, the images resulting from the “good” seed are more aligned. This emphasizes the dependency on started latent noise present at the beginning of the prompt, which remains a prominent factor of performance.
Figure 11: Generation of images with the same prompt and a “bad” (left) of a “good” (right) random seed. top: SD 2 “a drawing of a red lion and a blue moon, pop art, 4k, highly detailed” seed 27, 17. IF “a photo of a purple frog and a rainbow piano”, seed 19, 24.
5 Conclusion and Perspective
We proposed a new metric to automatically quantify the performances of T2I models in terms of prompt-image alignment. Contrary to previous research efforts, it is based on prompt templates, that allow a finer analysis with regard to the syntax of the prompt. Hence we showed that the alignment performances of most T2I models drop significantly with the number of objects specified in the prompt and that effect is even more critical for color attribution. Our metric also allows us to study the influence of the input seed at inference. We showed that there exist some seeds that systematically result in better output images than others, and that it generalizes on objects out of the set used to determine them. It draws possible future research toward the mining of such “good seeds”, as a complementary activity to prompt engineering to optimize the outputs of T2I models.
Acknowledgement
This work was granted access to the HPC resources of IDRIS under the allocation 2022-AD011014009 made by GENCI. This publication was also made possible by the use of the FactoryIA supercomputer, financially supported by the Ile-De-France Regional Council.
Supplementary Material
Appendix A Proof of Proposition 1 and its generalization
We give the proof of proposition 1, that is easy to establish, then we derive a formula giving the number of templates in a more general realistic case (proposition 3). For this, we introduce an intermediate step (proposition 2) and most importantly a set of notation that facilitates the derivation of the final analytic formula.
Let us consider a collection of object sets $\{\mathcal{O}_{i}\}_{i=1}^{N}$ and a collection of attribute sets $\{\mathcal{A}_{i}\}_{i=1}^{N}$ that are used to define a template with $N$ objects, each being qualified by an attribute. Hence, at position $i$ (without considering the context), the prompts results from the inference of the template and contains a named object $o_{i}\in\mathcal{O}_{i}$ that is qualified by a (color) adjective $a_{i}\in\mathcal{A}_{i}$.
Proposition 1:
if $|mO|\geq N$, $|\mathcal{A}|\geq N$ and $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$ and $\forall(i,j)\in[\![1,N]\!]^{2}$ s.t $i<j$, we force $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$, thus the number of unique prompts generated by the template is $\frac{|\mathcal{O}|!|\mathcal{A}|!}{\left(|\mathcal{O}|-N\right)!\left(|\mathcal{A}|-N\right)!}$
Proof.
Without attribute, each prompt contains $N$ objects that should be different thus the number of unique possible prompt is the number of (arrangements) $N$-permutations of $|\mathcal{O}|$ thus $\frac{|\mathcal{O}|!}{\left(|\mathcal{O}|-N\right)!}$. Similarly if one considers the attributes only that must be different at each position $i$ we have $N$-permutations of $|\mathcal{A}|$. Finally, since both are independent the final number of unique prompts generated by the template is the product of both. ∎
We can consider a generalization where the sets are different at each position, that is we remove the conditions $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$ from proposition 1. For example, such template could be defined to have a vehicle at first position, a fruit at the second position and an animal at the third one. The intersection of the sets at each position may also be non-empty, for example the vehicle may be $\{blue,red,green,black,yellow,white\}$, the fruit $\{red,green,yellow\}$ and the animal $\{green,black,yellow,white\}$. Similarly, the attributes can be repeated in the prompt e.g if one wants “a yellow car and a red apple and a red elephant”.
Proposition 2:
if $\forall i\in[\![1,N]\!]$, $|\mathcal{O}_{i}|\geq N$, $|\mathcal{A}_{i}|\geq N$ thus the number of unique prompts generated by the template is $\prod_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$. Hence, if $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$ thus the number of unique prompts generated by the template is $(|\mathcal{O}|.|\mathcal{A}|)^{N}$
Proof.
At any position $i$ any all the attributes $a_{i}\in\mathcal{A}_{i}$ can be associated with the object $o_{i}\in\mathcal{O}_{i}$ thus it gives $|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$ possibilities. Since the repetition are allowed, we have a structure of arborescence (rooted tree), thus the total number of prompts is the product at any position, thus $\prod_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$. ∎
However, if the sets of objects are not exclusive we consider that we should not allow any repetition, since it would be strange to require “a car and an apple and a car”. Such repetition can nevertheless be allowed if the object appears several times with different attributes. In other words, if $\forall i\in[\![1,N]\!]$, $|\mathcal{O}_{i}|\geq N$, $|\mathcal{A}_{i}|\geq N$ and $\forall(i,j)\in[\![1,N]\!]^{2}$ s.t $i<j$ we force $o_{i}\neq o_{j}$ if $a_{i}=a_{j}$.
Table 4: Example of objects with attributes, above dashed line: that can appear at a specific location and are thus in $\mathcal{U}_{i}^{0}$; below dashed line: that can appear at several locations in the prompt. One notes that in that case $\mathcal{O}_{1}=\{^{\prime}truck^{\prime},^{\prime}bike^{\prime},^{\prime}car^{\prime}\}$ and $\mathcal{A}_{1}=\{^{\prime}blue^{\prime},^{\prime}red^{\prime}\}$
$\mathcal{U}_{1}$ $\mathcal{U}_{2}$ $\mathcal{U}_{3}$
blue truck green banana green tiger
red truck purple banana blue tiger
blue bike green cherry green bear
red bike purple cherry blue bear
red car purple apple yellow apple
[…]
green apple green apple
Table 4: Example of objects with attributes, above dashed line: that can appear at a specific location and are thus in $\mathcal{U}_{i}^{0}$; below dashed line: that can appear at several locations in the prompt. One notes that in that case $\mathcal{O}_{1}=\{^{\prime}truck^{\prime},^{\prime}bike^{\prime},^{\prime}car^{\prime}\}$ and $\mathcal{A}_{1}=\{^{\prime}blue^{\prime},^{\prime}red^{\prime}\}$
configuration $\mathcal{P}$
pos. 1 pos. 2 pos. 3
blue car green apple $\mathcal{U}_{3}^{0}$$\{3\}$
blue car $\mathcal{U}_{2}^{0}$ green apple $\{2\}$
blue car $\mathcal{U}_{2}^{0}$ $\mathcal{U}_{3}^{0}$$\{2,3\}$
$\mathcal{U}_{1}^{0}$ green apple blue car $\{1\}$
$\mathcal{U}_{1}^{0}$ green apple $\mathcal{U}_{3}^{0}$$\{1,3\}$
$\mathcal{U}_{1}^{0}$ $\mathcal{U}_{2}^{0}$ blue car $\{1,2\}$
$\mathcal{U}_{1}^{0}$ $\mathcal{U}_{2}^{0}$ green apple $\{1,2\}$
Table 5: Illustration of the set $\mathcal{P}$ (fourth col.) for the example of Table 4. The notation $\mathcal{U}_{i}^{0}$ means that any colored object of $\mathcal{U}_{i}^{0}$ can be used at that position
In that case, things depend on the overlap between the combination of attributes and objects at each position. In other words, if the attributes are colors, it depends on how many times the same colored object can appear at different positions in the prompt. Let us consider the sets $\mathcal{U}_{i}=\mathcal{A}_{i}\times\mathcal{O}_{i}$ of objects with attributes at position $i$ in the prompt. Since we have $o_{i}\neq o_{j}$ if $a_{i}=a_{j}$, each $\mathcal{U}_{i}$ can be split into two disjoint subsets. The subset $\mathcal{U}_{i}^{0}$ contains the objects with attributes that can appear at position $i$ only, and its complement in $\mathcal{U}_{i}$ contains the objects with attributes that can appear at another position $j\neq i$ in the prompt.
The number of prompts that the template can generate depends on the overlap between these complements of the $\mathcal{U}_{i}^{0}$. For instance, the “blue car” can appear at positions 1 and 3, while the “green apple” can appear at positions 2 and 3 as illustrated in Tab. 4. The objects that can appear at several positions define some configurations that are uniquely defined by the set of positions where they appear in the prompt.
Symmetrically, these configurations are also defined by the sets of positions where it appears only object at a unique position (thus from $\mathcal{U}_{i}^{0}$). We note $\mathcal{P}$ the set of these last positions (See Tab. 5 for an illustration of these sets). Note that the same element several times can appear several times in $\mathcal{P}$, actually when several objects with attributes can appear at the same position (e.g both “blue car” and “green apple” can appear at position 3 in our example). Formally, $\mathcal{P}$ is a set that contains the sets of indexes of the non-empty positions of all the N-uples that verify $(\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0})\cap(\mathcal{U}_{j}\text{\textbackslash}\mathcal{U}_{j}^{0})\neq\emptyset$. Using this formalization of the problem, we have:
Proposition 3:
Let $\forall i\in[\![1,N]\!]$, $|\mathcal{O}_{i}|\geq N$, $|\mathcal{A}_{i}|\geq N$. If $\forall(i,j)\in[\![1,N]\!]^{2}$ s.t $i<j$ and $a_{i}=a_{j}$, we force $o_{i}\neq o_{j}$, thus the number of unique prompts generated by the template is:
$\prod_{i=1}^{N}|\mathcal{U}_{i}^{0}|+\sum_{P\in\mathcal{P}}\prod_{i\in P}|\mathcal{U}_{i}^{0}|$ (4)
Proof.
Since the elements of the $\mathcal{U}_{i}^{0}$ are different and appear at a unique position, they generate $\prod_{i=1}^{N}|\mathcal{U}_{i}^{0}|$ templates.
For any configuration in $\mathcal{P}$, an object with attributes (from one of the $\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0}$) appear at most once since if $a_{i}=a_{j}$ thus $o_{i}\neq o_{j}$.
For each configuration $P\in\mathcal{P}$ (such as the lines of Tab. 5) the number of templates generated is the product of all the $|\mathcal{U}_{i}^{0}|$ for that configuration, thus $\prod_{i\in P}|\mathcal{U}_{i}^{0}|$ for each configuration of position $P$. If we sum for all the possible $P$ in $\mathcal{P}$, it results in the number of templates generated with objects with attributes that can appear at several positions.
The sum of both terms gives the value in Equation 4
Hence the exact number of prompts generated depends on the possible overlaps between objects with attributes, both the number of element and the place the can appear or not. By introducing more notations on these elements and their position, one could derive a formula but it would be tedious, without obvious interest in practice. Indeed, $\mathcal{P}$ can be built easily through a tree structure, by considering iteratively all the sets $(\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0})\cup\{\emptyset\}$, and allowing only the nodes that did not previously appear and the branch corresponding to non-empty set in $\mathcal{P}$ (such as in Tab. 5).
Appendix B Prompt Templates
We detail the templates and the COCO labels used in the study.
B.1 Without attribute
The 24 COCO labels used for the general study of sections 4.1 and 4.2 are the following: bicycle, car, motorcycle, truck, fire hydrant, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, banana, apple, broccoli, carrot, chair, couch, oven, refrigerator. The prompt template is “a photo of $det(o_{1})$ $o_{1}$ and $det(o_{2})$ $o_{2}$”.
For the semantic studies, end of section 4.3, we use the vehicles, animals, and food labels i.e. bicycle, car, motorcycle, airplane, bus, train, truck, boat, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake. The template used is “a photo of $det(o_{1})$ $o_{1}$ and $det(o_{2})$ $o_{2}$”.
To evaluate the capacity of models to represent multiple objects, we use the following template :
  • 1 objects : “a photo of $det(o_{1})$ $o_{1}$
  • 2 objects : “a photo of $det(o_{1})$ $o_{1}$ and $det(o_{2})$ $o_{2}$
  • 3 objects : “a photo of $det(o_{1})$ $o_{1}$” next to $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$
  • 4 objects : “a photo of $det(o_{1})$ $o_{1}$” next to $det(o_{3})$ $o_{2}$ with $det(o_{3})$ $o_{3}$ and $det(o_{4})$ $o_{4}$
We do not use commas because it tends to reduce the score. We report in Tab 6 the comparison between the prompt “a photo of $det(o_{1})$ $o_{1}$ next to $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$” and the prompt with a comma “a photo of $det(o_{1})$ $o_{1}$, $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$”.
Model w comma w/o comma
IF 0.12 0.21 ${}_{(+0.09)}$
SD 1.4 0.01 0.02 ${}_{(+0.01)}$
SD 1.4 A&E 0.13 0.17 ${}_{(+0.04)}$
SD 2 0.13 0.15 ${}_{(+0.02)}$
SD 2 A&E 0.17 0.21 ${}_{(+0.04)}$
unCLIP 0.13 0.10 ${}_{(-0.03)}$
Table 6: Comparison of the TIAM for two templates with 3 entities, separated by commas or related with words.
B.2 With Attribute
To evaluate the attribute binding in section 4.4, we used the following templates:
  • one object : “a photo of $det(a_{1})$ $a_{1}$ $o_{1}$
  • two objects : “a photo of $det(a_{1})$ $a_{1}$ $o_{1}$ and $det(a_{2})$ $a_{2}$ $o_{2}$
[…]
Appendix C Reference colors
We plot on Fig 13 the interpolation of the best examples in Lab on the CIE 1931 Chromaticity Diagram.
Figure 12: L2 norm distance between our reference colors.
We do not use our reference colors orange and brown because they are respectively too close to red and black in the CIELab space (Fig. 12).
Figure 13: CIE 1931 Chromaticity Diagram with the best example for each color. When multiple best examples for one color, we compute one best example by averaging the value in the CIELAB space.
Appendix D Occurrence of objects on images
We show in our experiments (section 4.3) that the initial objects in the template tend to appear more frequently than objects inserted subsequently and reinforce the observation that the concept that is expressed earlier in the prompt has more chances to appear in the final image.
We present the result for two objects in Fig. 14 and for three objects in Fig. 15.
Figure 14: The proportion of occurrences of each object, based on its positions in the prompt. The template of the prompt includes two objects.
Figure 15: The proportion of occurrences of each object, based on its positions in the prompt. The template of the prompt includes three objects.
Appendix E Determining the minimum number of images to generate
In Fig. 16 we report the TIAM score as a function of the number of generated images per prompt. The score stabilizes from 16 images. We chose to compute with 32 images to ensure robustness.
Figure 16: TIAM as the function of the number of generated images per prompt.
Appendix F Detection/Segmentation details
We use the largest YOLOv8 for segmentation [^6] https://github.com/ultralytics/assets/releases/download/v0.0.0/yolov8x-seg.pt. During segmentation inference, we set up the object confidence threshold for detection to 0.25 and the intersection over union IoU threshold for NMS to 0.8. We compute the score with different confidence thresholds and observe that the score decreases linearly as the threshold values increase (Fig. 17).
Figure 17: TIAM as a function of the YOLO object confidence threshold.
Appendix G Text-to-image models Setup
All generations of images were done on Nvidia A100 SXM4 80 Go using float16. We list the main parameters for the different models used. If not mentioned, we use the default parameters from the library diffusers [30] (version 0.16.1).
SD 1.4  [^7] https://huggingface.co/CompVis/stable-diffusion-v1-4 and SD 2  [^8] https://huggingface.co/stabilityai/stable-diffusion-2-base
The models produce images of size 512$\times$512.
unCLIP  [^10] https://huggingface.co/kakaobrain/karlo-v1-alpha
The model produces images of size 256$\times$256.
  • 25 inference steps for the prior, 25 inference steps for the decoder, and 7 steps of Super-resolution
  • Prior guidance scale: 4, decoder guidance scale: 8
  • Scheduler is the UnCLIPScheduler, a modified DDPM scheduler designed for this model.
IF
It exists a different configuration of the Deepfloyd IF. We use for the first stage the L version [^11] https://huggingface.co/DeepFloyd/IF-I-L-v1.0 with 100 inference steps and the second stage the M version [^12] https://huggingface.co/DeepFloyd/IF-II-M-v1.0 with 50 inference steps. We use both the DDPMS scheduler and guidance scale of 7 for the first stage and 4 for the second stage. With this configuration, we produce images of size 256$\times$256.
Appendix H Semantic Link
To investigate the impact of semantic relationships between objects, we select 28 COCO labels from three macro-classes, vehicles, animals, and foods, and generate images using a template with 2 objects. In order to study the influence of semantic links between objects with the same prompt, we consider the following dissimilarity metric on the set $\mathcal{O}$.
$\forall o_{x},o_{y}\in\mathcal{O},o_{x}\neq o_{y},\\ \hskip 28.45274ptd(o_{x},o_{y})=d(o_{y},o_{x})=\frac{\text{TIAM}_{z_{i}}+\text{TIAM}_{z_{j}}}{2}$ (5)
$\forall o_{x}\in\mathcal{O},\hskip 28.45274ptd(o_{x},o_{x})=0$ (6)
[…]
Figure 18: MDS on the objects score dissimilarity for SD 1.4.
For all the models we observe some clusters of objects from the same macro-class, in particular animals. It shows that when two objects are semantically close, they tend to be harder to be generated in the same image. Tang et al. [27] obtain results in the same vein, showing that it is easier to generate a non-cohyponym than a cohyponym.
However, the effect remains slight, suggesting that the cohyponymy has either an indirect or minor link to this difficulty of generation. We quantified the effect by computing the correlation between the TIAM score for all the templates with two objects and the semantic distance between the two objects. We used various methods to estimate the semantic distance, including Wu-Palmer, the CLIPscore, and the cosine similarity between the embedding of the token in the prompt (before the attention of the transformer) for SD 1.4 and SD 2. We report the results in Tab 7. For all distances, these correlations are negative (confirming the effect) but their absolute values are less than 0.3 (confirming the effect is slight).
Correlation with SD 1.4 SD 2
CLIPscore -0.09 -0.1
Embedding matrix -0.06 -0.12
Wu-Palmer -0.28 -0.28
Similarity in context -0.17 -0.24
Key (mean±$\pm\sigma$) $-0.17\pm{0.08}$ $-0.17\pm{0.09}$
Value (mean±$\pm\sigma$) $-0.19\pm{0.07}$ $-0.17\pm{0.08}$
Table 7: Correlation between the semantic distance between two objects ($o_{1}$ and $o_{2}$) and the alignment of the image generated by Stable Diffusion TIAM and a template including $o_{1}$ and $o_{2}$. The objects belong to three macro-classes (vehicles, animals, foods). The CLIPscore is the cosine similarity between the CLIP textual features (first eot embedding) of $o_{1}$ and $o_{2}$. Embedding matrix is the cosine similarity between the representation of tokens $o_{1}$ and $o_{2}$ in the semantic space prior to attention in the CLIP model. Wu-palmer stands for the Wu-Palmer similarity between $o_{1}$ and $o_{2}$. The similarity in context is the cosine similarity between the representation of the tokens $o_{1}$ and $o_{2}$ in the last hidden states of the embedding of the prompt by the CLIP textual encoder. Finally, Key and Value indicate that we compute the average score and standard deviation ($\sigma$) of all the correlations of the representations Key and Value. We calculated these correlations for each attention head of each layer of the U-net.
Figure 19: MDS on the objects score dissimilarity for SD 2.
Figure 20: MDS on the objects score dissimilarity for unCLIP.
Figure 21: MDS on the objects score dissimilarity for IF.
Appendix I Attribute binding
Figure 22: TIAM per object i.e. proportion of correct generated object with the correct binding.
We report the TIAM per object on Fig. 22, showing that the first object is more often generated and correctly colored. We compute the binding success rate, but by differentiating by colors for attributes in the first position (Fig. 23) and attributes in the second position (Fig. 24). We observed that the models face greater difficulty in assigning green and blue colors when two objects are involved (parallel with a single label case). It is worth noting that IF performs better than other models.
Figure 23: Binding success rate for the first object, among the first objects correctly detected.
Figure 24: Binding success rate for the second object, among the second objects correctly detected.
Appendix J Latent Diffusion Model
We remind the architecture of the Latent Diffusion Model of Rombach et al. (2021) in Figure 26, since the theoretical explanation of Section 4.2 of the main paper relies on it.
Figure 25: Standardized score per seed for SD 1.4 and SD 2. We observe that both of them have globally the same “good” and “bad” seeds (see subsection 4.2 for explanations).
Figure 26: Architecture of the Latent Diffusion Model [24].
During the investigation of the seed performances of the models, we made a noteworthy finding. We observed that the SD models exhibited similar behavior in relation to the seeds (if we do not consider the performance gap between the score of the model) i.e. when we standardize the score of each seed for each model (Fig 25) the score exhibits a remarkable similar trend for both models with the same “good” and “bad” seeds. However, we explain in the article (section 4.2) that the “good” and “bad” seeds are specific to each model, which may seem contradictory at first glance.
To explain this point, we need to remind the training of diffusion models. Let $x_{0}$ be an original image. We had on the image a scaled quantity of noise $\epsilon\sim\mathcal{N}(0,1)$ to obtain $x_{t}$ ($\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$) and the U-net $\epsilon_{\theta}$ try to predict with the added noise (the loss function ($\mathbb{E}_{\mathcal{E}(I),\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}\right]$). At each training step a $t$ is drawn the model must predict the noise. In the case of LDM, just replace the $x$ with $z$ because the diffusion process is in the latent space.
As the two models are trained on the same data, we hypothesize that they are trained on the same $z_{t}$ i.e. they are trained to predict the same $\epsilon$. However $\epsilon$ is random, we suspect that all the $z_{t}$ are precomputed and the associate $\epsilon$ saved, and they used the same for the training of both models. Because of this particularity during the training, the models try to retrieve the $z_{0}$ from multiple possible $t$. The models learn a similar path of reverse diffusion. At inference, when we draw a similar noise $\chi_{T}$, the reverse process will be similar and conduct to near $\chi_{t}$ and finally to a near $\chi_{0}$. We show in Fig.8 that with the same prompt and the same starting noise we exhibit strong composition similarity that explains the parallel performance of the seed.
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 8: We present images obtained with the same seed and same prompt for SD 1.4 (left) and SD 2 (right). Note how similar the compositions are. The prompt and seed are respectively ”a photo of a blue circle” seed 17, ”a photo of a cat” seed 27, ”a photo of a pink elephant” seed 18, and ”a photo of a penguin” seed 45.